For n=47 and X=43, find the exact proportions p and q. = 0 :
Answer #1

n = 47 and x = 43

pcap = 43/47 = 0.914893617

qcap = (47 - 43)/47 = 4/47 = 0.085106383

If answer is required in fraction use fractions